HETAV PATEL

Grade 11,

When a capillary tube is immersed vertically in water the capillary rise is 3 cm. if the same capillary tube is inclined at angle of 60° to the vertical, the length of the water column in the capillary tube above that of the outside level is (a) 6 cm (b) 1 cm (c) 8 cm (d) Zero

When a capillary tube is immersed vertically in water the capillary rise is 3 cm. if the same capillary tube is inclined at angle of 60° to the vertical, the length of the water column in the capillary tube above that of the outside level is
(a) 6 cm (b) 1 cm
(c) 8 cm (d) Zero

Grade:Upto college level

2 Answers

Saurabh Koranglekar
3 years ago

Vertical Capillary rise will be constant
l cos60=h1​
l=2×3cm=6cm
Vikas TU
3 years ago
Vertical height of the liquid column h(=3cm) in inclined capillary tube will be the same as that in the vertical capillary tube when immersed in a liquid. If l is the length of liquid column along the tube in the inclined tube, then
cosθ=h/l
or
l=h/cosθ
=3/cosθ60∘
=6cm.